evaluate the following integral
∫ e^(x) + 1 / e^(x) + x dx

Grade:12

2 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{e^{x}+1}{e^{x}+x}dx
e^{x}+x = t
(e^{x}+1)dx = dt
I = \int \frac{1}{t}dt
I = log(t) + c
Jitender Singh IIT Delhi
9 years ago
Ans:
Hello student,
Please find the answer to your question below

I = \int \frac{e^{x}+1}{e^{x}+x}dx
e^{x}+x = t
(e^{x}+1)dx = dt
I = \int \frac{1}{t}dt
I = log(t) + c
I = log(e^{x}+1)+c